If \\(a=3i-2j+k\\), \\(b=2i-4j-3k\\) and \\(c=-i+2j+2k\\), find the magnitude of \\(a+b+c\\)
1
Expert's answer
2020-04-16T18:32:09-0400

"\\a=3i-2j+k\\ \\b=2i-4j-3k\\ \\c=-i+2j+2k\\a+b+c=4i-4j"

its magnitude will be "\sqrt{4^2+4^2}=4\sqrt2"